ganeshie8 (ganeshie8):

if nothing can go beyond the speed of light, then how a BUFFER can have a delay as low as 50 picosecond or how ICs can work at Ghz range (> 10^9 cycles per second ) ? i thought i figured this out some time back... but right now im blank again... . can somebody pls enlighten me

OpenStudy (ghazi):

see basically speed of electromagnetic is C=lambda* frequency ..now when you talk about frequency of GHz you must see wavelength is decreased proportionally therefore product of frequency and wavelength will give speed of light again...hence you're done

OpenStudy (ghazi):

i mean electromagnetic wave

ganeshie8 (ganeshie8):

i understand that a bit.. wavelength * frequency = C im not able to relate wavelength to ICs. some ICs operate at 10Ghz. they turn on and off 10 billion times per second... how any event can happen faster than speed of light is what i dont understand... (sorry if it sounds very primitive/stupid q... i have only opened my books now n getting all weird questions... )

OpenStudy (anonymous):

ohh.........., it's an interesting question.

OpenStudy (shane_b):

What @ghazi said answers the question. You can have extremely high frequencies if you also have very small wavelengths. Using your example: \[\frac{c}{f}=wavelength\]\[\frac{2.998x10^8m/s}{1.000x10^{10}Hz}=0.02998m\]

ganeshie8 (ganeshie8):

yeah i see it. IC is very small in size. so the wave travels very small distance few mm at max. i must be stupid. i get it thanks a lot :D

OpenStudy (ghazi):

hope it's clear now...by the way good question to think